Dear Sir/Ma'am,


Trust you are doing well.


Can you help me to slove below question please.

Your swift action would be highly appreciated.


Q. The demand for dog treats is represented by the following equation.


QD =300-50P Qs = - 100 + 150P


Determine equilibrium Quantity Demand, Quantity Supply and Price.


1
Expert's answer
2021-11-03T17:49:03-0400

Solution:

At equilibrium: QD = QS

300 – 50P = -100 + 150P

300 + 100 = 150P + 50P

400 = 200P

P = 2

Equilibrium price = 2


Substitute in the demand or supply function to derive equilibrium quantity:

QD = 300 – 50P = 300 – 50(2) = 300 – 100 = 200

QS = -100 + 150P = -100 + 150(2) = -100 + 300 = 200

Equilibrium quantity = 200